# Vlaser: Vision-Language-Action Model with Synergistic Embodied Reasoning

### v1

@@ A.2 Data Generation Details
In-Domain Data for downstream VLAs We generate 2 million in-domain multimodal data samples to facilitate direct transfer learning for downstream Vision-Language Agents (VLAs) during fine-tuning. These data are collected from two distinct robotic platforms: the WidowX Robot (Walke et al., 2023a) and the Google Robot (Brohan et al., 2023b; a) within the SimplerEnv (Li et al., 2024c) simulation framework. The dataset mainly encompasses three systematically designed question-answer types: 1) General QA, which queries the robot’s current state and requests next few action plans; 2) Grounding QA, which requires the robot to localize points and bounding boxes as the actionable affordances; 3) Spatial Reasoning QA, which probes relative spatial relationships and geometric properties of objects in the scene. Detailed prompt templates and representative examples for each data category are provided in Figure 3 (General QA), Figure 4 (Grounding QA), and Figure 5 (Spatial Reasoning QA), respectively. We use Qwen2.5VL-7B (Bai et al., 2025) as the base model to generate the above data items.
@@ A.4 Qualitative Demonstration
Qualitative Samples for Embodied Reasoning. To gain in-depth insights into Vlaser, we provide extensive visualizations on embodied reasoning tasks including Embodied QA 7, grounding 8, spatial intelligence 9 and planning tasks in Simulation 10. These examples not only validate the generalizability of Vlaser, but also confirm the superior capabilities of Vlaser across all tasks.

### v2

@@ 6 Conclusion and Discussion
ETHICS STATEMENT
This research adheres to the ICLR 2026 ethical guidelines and upholds the principles of responsible research. We ensure that no personally identifiable, sensitive, or harmful data were used. Our experiments were based on publicly available datasets and did not involve any human subjects or vulnerable groups. We have considered the potential societal impact of our methods, including the risk of misuse, and believe that these contributions primarily advance scientific understanding and do not pose foreseeable harm.
REPRODUCIBILITY STATEMENT
We follow the reproducibility guidelines in the ICLR 2026 author guidelines. We will open source code, configuration files, and scripts to reproduce our results, including dataset construction, model training, and evaluation, on platforms such as GitHub and Huggingface as soon as possible.
@@ Appendix A Appendix
THE USE OF LARGE LANGUAGE MODELS (LLMS)
We used a large language model, ChatGPT (OpenAI, 2025), only for grammar check and correction. All the writings was manually reviewed by the authors. Crucially, AI was not used to generate any research data, statistical analyses, figures, or conclusions. The authors take full responsibility for the final content and any errors herein.
@@ A.2 Data Generation Details
In-Domain Data for downstream VLAs We generate 2 million in-domain multimodal data samples to facilitate direct transfer learning for downstream Vision-Language Agents (VLAs) during fine-tuning. These data are collected from two distinct simulation platforms: the WidowX Robot (Walke et al., 2023a) and the Google Robot (Brohan et al., 2023b; a) within the SimplerEnv (Li et al., 2024d), as well as the Aloha-AgileX Robot from RoboTwin2.0 (Chen et al., 2025c). The dataset mainly encompasses three systematically designed question-answer types: 1) General QA, which queries the robot’s current state and requests next few action plans; 2) Grounding QA, which requires the robot to localize points and bounding boxes as the actionable affordances; 3) Spatial Reasoning QA, which probes relative spatial relationships and geometric properties of objects in the scene. Detailed prompt templates and representative examples for each data category are provided in Figure 3 (General QA), Figure 4 (Grounding QA), and Figure 5 (Spatial Reasoning QA), respectively. We use Qwen2.5VL-7B (Bai et al., 2025) as the base model to generate the above data items.
To further enhance the quality and diversity of the generated in-domain QA data, we implement a post-processing data filtering pipeline. In line with established practices in dataset construction (Wang et al., 2023; Cui et al., 2023; Li et al., 2024b), we employ an LLM-as-a-judge approach (Li et al., 2025) to score each generated data sample on a scale of 1 to 3, using Qwen2.5VL-32B (Bai et al., 2025) as the judge model. The detailed instruction prompts for General QA, Spatial QA, and Grounding QA are provided below. We filter out all samples assigned a score of 1, which account for approximately 10% of the initially generated data. This outcome reflects the overall high quality and diversity of our in-domain dataset. To further validate the reliability of the LLM-based evaluation, we randomly selected a subset of the scored data (covering all score levels from 1 to 3) for human reassessment by three human experts. The results show that the LLM judge’s scores align with human ratings in nearly 80% of cases, confirming the consistency and reliability of our automated evaluation and data filtering pipeline.
@@ A.4 Qualitative Demonstration
Qualitative Samples for Embodied Reasoning. To gain in-depth insights into Vlaser, we provide extensive visualizations on embodied reasoning tasks including Embodied QA 9, grounding 10, spatial intelligence 11 and planning tasks in Simulation 12. These examples not only validate the generalizability of Vlaser, but also confirm the superior capabilities of Vlaser across all tasks.